OpenStudy (anonymous):

y-intercept of -2 and slope of 1/3

OpenStudy (skullpatrol):

What are you asking for?

OpenStudy (anonymous):

equation of lines

OpenStudy (skullpatrol):

How many lines?

OpenStudy (anonymous):

Slope intercept form will do the trick. It ha exactly what it sounds like, the slope,and the y intercept. It loooks like: y = mx + b m is the slope, and b is the y intercept just plug it in! ex: y int: -5 Slope: 1/2 Equation of the line: y = 1/2x -5 Slope y int

OpenStudy (anonymous):

but how youput it on the graph????

OpenStudy (skullpatrol):

That was not in your question!!!!

OpenStudy (anonymous):

You want to graph it now??

OpenStudy (anonymous):

please

OpenStudy (anonymous):

Ok..

OpenStudy (anonymous):



OpenStudy (anonymous):